### v1

# FOSP: Fine-tuning Offline Safe Policy through World Models

Abstract
Model-based Reinforcement Learning (RL) has shown its high training efficiency and capability of handling high-dimensional tasks. Regarding safety issues, safe model-based RL can achieve nearly zero-cost performance and effectively manage the trade-off between performance and safety. Nevertheless, prior works still pose safety challenges due to the online exploration in real-world deployment. To address this, some offline RL methods have emerged as solutions, which learn from a static dataset in a safe way by avoiding interactions with the environment. In this paper, we aim to further enhance safety during the deployment stage for vision-based robotic tasks by fine-tuning an offline-trained policy. We incorporate in-sample optimization, model-based policy expansion, and reachability guidance to construct a safe offline-to-online framework. Moreover, our method proves to improve the generalization of offline policy in unseen safety-constrained scenarios. Finally, the efficiency of our method is validated on simulation benchmarks with five vision-only tasks and a real robot by solving some deployment problems using limited data.
Keywords: Safe Model-based Reinforcement Learning, Offline-to-Online
1 Introduction
Model-based Reinforcement Learning (MBRL) has proven to be a powerful tool to enhance sample efficiency by constructing a world model of the environment[1, 2, 3, 4, 5]. The world model reduces random explorations by predicting the outcomes of different actions, leading to a fast convergence speed during online fine-tuning after minimal interactions [6]. Moreover, it significantly improves action safety in high-dimensional tasks such as vision-based operations due to its prediction ability[7, 8]. Consequently, We consider utilizing model-based methods to tackle safety-critical problems in real-world deployment.
Offline RL has been widely studied within the academic community, using pre-collected datasets to learn policies without exploration[9]. At the offline training stage, safety constraints will be incorporated into policy learning, which means learning to be safe other than learning safely. In this way, offline trained policy is promising to avoid constraint violations during the online deployment. However, offline dataset cannot fully simulate complicated real-world environments and will cause out-of-distribution (OOD) issues in the case of unseen data. Such generalization tasks are significant challenges for offline learning.
In this work, we propose developing an offline-to-online approach to address the generalization issue with safety guarantees. We consider the problem of pretraining an offline policy and a world model, and then fine-tuning them safely during online interactions. However, such a strategy usually causes degradation of policy performance and a lack of safety after fine-tuning[10]. The main reason lies in too many mixed constraints (both soft and hard), such as safety constraints and behavior regularization, making it difficult to ensure both optimality and safety. Meanwhile, the world model will introduce prediction errors due to distribution shift during the fine-tuning phase, leading to increasing safety risks. Therefore, we seek to unify and handle mixed constraints, and use the offline safe policy as a guidance for online correction. Based on this key insight, we proposed FOSP (Fine-tuning Offline Safe Policy through World Models), a method that enhances offline safe training and bridges model-based offline training with online fine-tuning.
To evaluate our method, we design different experiments in the simulation including offline training and online fine-tuning. It shows robust performance in both offline and online fine-tuning tasks in Safety-Gymnasium[11] and achieves nearly zero-cost. We also deploy FOSP in a real-world Franka manipulator for trajectory planning tasks. It turns out that our method performs well across tasks with different safety regions and can be transferred to unseen scenarios through few-shot fine-tuning.
We summarize our contributions as follows: (1) To the best of our knowledge, FOSP is the first approach that extends model-based RL to safe offline-to-online areas. (2) It handles the trade-off between the performance and constraints satisfaction across offline and online training phases on various vision-only tasks. (3) In real-world deployment, it does not need sim-to-real transfer and can be safely fine-tuned in unseen safety-constrained scenarios. (4) To inspire further research in this area, we will publicly release our code and dataset.
Figure 1: Fine-tuning offline safe policy through world models. We propose an approach to offline pretrain and online fine-tuning the world model. We also consider the safe constraints and generalize it to unseen safe regions.
2 Related Works
Offline-to-online RL
Offline RL, which trains a policy by leveraging a large amount of existing data, is valuable to real-world scenarios[12, 13]. Prior offline RL works focus on addressing the distribution shift and value overestimation problem[14, 15, 16, 17, 18, 19]. Recently, in-sample methods have demonstrated their robust performance by avoiding out-of-distribution (OOD) state-action pairs[20, 21, 22, 23, 10, 24]. Unlike training solely on offline data, offline-to-online RL involves further fine-tuning the offline policy online. And many offline RL algorithms have extended to online fine-tuning[20, 10]. But these methods have shown that the more effective the offline RL method is, the worse online fine-tuning performance is[25]. Some previous works have explored how to effectively fine-tune from offline pretrained policy, including balanced sampling in replay buffer[26], policy calibrating[27, 28], and parameter transferring[29, 30]. However, existing solutions are restricted by policy transitions or the need to estimate the density of the online policy distribution[31]. PEX[32] is an approach that bridges offline and online algorithms and is adaptive for fine-tuning different online methods with offline policy guidance. As it is shown as an effective approach in offline-to-online training, we extend it to model-based algorithms that not only require fine-tuning the policy but also the model.
Safe RL
Safe RL focuses on optimizing objectives with different constraints[33]. Many methods are based on policy search algorithms that ensure adherence to near-constraints iteratively[34] and Lagrangian methods that convert it to an unconstrained problem to solve[35, 36, 37]. Model-based RL methods utilize world models for action selection so that they ensure safety by background planning. Some recent works are trying to address challenges remaining in vision-input tasks[38, 39, 3]. LAMDBA[7] is a method extending DreamerV1[1] to safe RL by using augmented Lagrangian. Safe-SLAC[40] incorporating the Lagrangian method into SLAC[41] improves the computation complexity of LAMBDA. SafeDreamer[8] is the latest work in safe model-based reinforcement learning. It enhances safety in online or background planning with DreamerV3[5], achieving state-of-the-art (SOTA) performance. However, prior works only aim to address issues in the online setting, which cannot avoid costs incurring during exploration in the real world. By contrast, some safe offline RL can learn policies from offline datasets by using OOD action detection[42, 43, 44], DICE-based theory[45], and sequence modeling[46, 47]. Nonetheless, they only utilize soft constraints which sometimes can not effectively decrease constraint violations. Some works consider the hard constraints by using control theory methods such as control barrier function (CBF)[48] and HJ reachability[49, 50, 51, 52]. Our work draws from these works but unifies all the constraints with model-based background planning.
3 Preliminaries
World Model
Our work extends DreamerV3[5], a powerful and robust baseline model-based RL using world model and actor-critic. It has the ability to model high-dimensional observations of the environment and increase the data efficiency of RL. The world model learns compressed image input representations and predicts future states and rewards sequences for decision-making. It is implemented as a Recurrent State-Space Model (RSSM)[3]. First, the model maps the high-dimensional observations $x_{t}$ to stochastic states $z_{t}$. And deterministic latent states $h_{t}$ are predicted by $z_{t}$ and actions $a_{t}$. By concatenating both of them, we get the model state $s_{t}=\{h_{t},z_{t}\}$. Finally, we facilitate $s_{t}$ to predict the next observations and rewards. Due to the safety considered in the environment, we add a cost decoder to the original model. All components in the world model are as follows:
latent $\begin{split}\begin{array}[]{ll}\text{latent encoder:}&{\bm{z}}_{t}\sim q_{\theta}({\bm{z}}_{t}\mid{\bm{h}}_{t},{\bm{x}}_{t})\\ \text{deterministic state:}&{\bm{h}}_{t}=f_{\theta}({\bm{z}}_{t-1},{\bm{h}}_{t-1},{\bm{a}}_{t-1})\\ \text{stochastic state:}&\hat{{\bm{z}}}_{t}\sim p_{\theta}({\bm{z}}_{t}\mid{\bm{h}}_{t})\\ \end{array}\end{split}\;\begin{split}\begin{array}[]{ll}\text{observation decoder:}&\hat{{\bm{x}}}_{t}\sim p_{\theta}({\bm{x}}_{t}\mid{\bm{z}}_{t},{\bm{h}}_{t})\\ \text{reward decoder:}&\hat{{\bm{r}}}_{t}\sim p_{\theta}({\bm{r}}_{t}\mid{\bm{z}}_{t},{\bm{h}}_{t})\\ \text{cost decoder:}&\hat{{\bm{c}}}_{t}\sim p_{\theta}({\bm{c}}_{t}\mid{\bm{z}}_{t},{\bm{h}}_{t})\\ \end{array}\end{split}$
The model is trained end-to-end by the ELBO or variational free energy of a hidden Markov model loss function with KL balancing:
$\mathcal{L}^{\text{model}}(\theta)=\operatorname*{\mathbb{E}}_{\tau\sim\mathcal{D}}\Big{[}\sum_{t=1}^{T}-\ln p_{\theta}({\bm{x}}_{t}\mid{\bm{s}}_{t})-\ln p_{\theta}({\bm{r}}_{t}\mid{\bm{s}}_{t})-\ln p_{\theta}({\bm{c}}_{t}\mid{\bm{s}}_{t})+$
$\mathbb{D}_{KL}[sg(q_{\theta}({\bm{z}}_{t}|{\bm{h}}_{t},{\bm{x}}_{t}))||p_{\theta}({\bm{z}}_{t}|{\bm{h}}_{t})]+\beta\mathbb{D}_{KL}[q_{\theta}({\bm{z}}_{t}|{\bm{h}}_{t},{\bm{x}}_{t})||sg(p_{\theta}({\bm{z}}_{t}|{\bm{h}}_{t}))]\Big{]}.$ (1)
The stop-gradient operator makes the representations more predictable in the imagination training.
Safe Model-based RL
Safe RL is frequently formulated as a Constrained Markov Decision Process (CMDP) $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathbb{P},\mathcal{R},\mathcal{C},\mu,\gamma)$[33]. $\mathcal{S}$ and $\mathcal{A}$ denote the state and action space. $\mathbb{P}(s^{\prime}|s,a)$ represents the transition probability from $s^{\prime}$ to $s$ under action $a$. $\mathcal{R}$ is reward space that maps $\mathcal{S}\times\mathcal{A}$ to $\mathbb{R}$. And $\mathcal{C}$ is a cost function set containing cost functions $c:\mathcal{S}\times\mathcal{A}\rightarrow[0,C_{max}]$. $\mu(\cdot):\mathcal{S}\rightarrow[0,1]$ is the initial state distribution and $\gamma\in(0,1)$ is the discount factor. Given a policy distribution $\pi_{\psi}$, we define the performance and constraints as follows:
$J^{\mathcal{R}}(\pi_{\psi})=\operatorname*{\mathbb{E}}_{{\bm{a}}_{t}\sim\pi,{\bm{s}}_{t+1}\sim p_{\theta},{\bm{s}}_{0}\sim\rho}\left[\sum_{t=0}^{T}r_{t}\big{\rvert}{\bm{s}}_{0}\right],$ (2)
$J^{\mathcal{C}}_{i}(\pi_{\psi})=\operatorname*{\mathbb{E}}_{{\bm{a}}_{t}\sim\pi,{\bm{s}}_{t+1}\sim p_{\theta},{\bm{s}}_{0}\sim\rho}\left[\sum_{t=0}^{T}c_{t}^{i}\big{\rvert}{\bm{s}}_{0}\right]\leq d^{i},\;\forall i\in\{1,\dots,C\},$ (3)
where $i$ denotes different safety constraints we want to avoid and $d^{i}$ are cost thresholds. Therefore, our safe model-based RL problem can be formulated as follows:
$\max_{\psi}J^{\mathcal{R}}(\pi_{\psi})\;\;\;\text{s.t. }\;J^{\mathcal{C}}_{i}(\pi_{\psi})\leq d^{i},\;\forall i\in\{1,\dots,C\}.$ (4)
4 Methods
In this section, we introduce our framework (Figure 1) for fine-tuning safe actions with offline pretrained world models. It alleviates the distribution shift error during offline training and balances the trade-off between performance and violations via the reachability estimation function. In the online fine-tuning phase, using offline-trained policy as guidance helps mitigate the performance drop and further improves the performance of the algorithm. In section 4.1, to address the distribution shift in offline training, we use the in-sample optimization approach to train critics. Next, we introduce an offline-to-online approach called model-based policy expansion to further improve online performance in section 4.2. The reachability guidance for safety consideration is provided in section 4.3.
4.1 In-sample Optimization for Offline Training
Value overestimation is an extensive challenge in offline training due to the distribution shift between the state-action pair from the dataset and generated by the learned policy. Moreover, model-based offline RL algorithms not only learn a policy but also a dynamic model $p_{\theta}$ by the offline dataset. It makes this challenge more severe that the latent dynamic estimation and return prediction are also affected. Therefore, we introduce the in-sample optimization to address this issue.
First, we consider a sequence of data $\mathcal{B}=({\bm{x}}_{1:T},{\bm{a}}_{1:T},{\bm{r}}_{1:T},{\bm{c}}_{1:T})$ sampled from the offline dataset. Then, the model generates latent states ${\bm{s}}_{1:T}\sim q_{\theta}({\bm{s}}_{1:T}|{\bm{h}}_{1:T},{\bm{x}}_{1:T})$ and uses them as initial states to generate rollouts:
$\hat{{\bm{a}}}_{1:T}^{h}\sim\pi_{\psi}({\bm{a}}|\hat{{\bm{s}}}_{1:T}^{h-1}),\ \hat{{\bm{s}}}_{1:T}^{h+1}\sim p_{\theta}({\bm{s}}|\hat{{\bm{a}}}_{1:T}^{h},\hat{{\bm{s}}}_{1:T}^{h}),\ \hat{{\bm{r}}}_{1:T}^{h}\sim p_{\theta}({\bm{r}}|\hat{{\bm{s}}}_{1:T}^{h}),\ \hat{{\bm{c}}}_{1:T}^{h}\sim p_{\theta}({\bm{c}}|\hat{{\bm{s}}}_{1:T}^{h}).$ (5)
We use $Q_{\phi}({\bm{s}},{\bm{a}})$ as the critic network and $\hat{Q}_{\phi}(s,a)$ as the target network. Similar to Implicit Q-learning (IQL)[24], we only use in-sample actions as calculated components of $\lambda$-return. Specifically, a large target value estimated by $Q_{\phi}$ might not represent an existing action leading to a good state due to the out-of-sample actions $\pi({\bm{s}})$. Thus, a separate value network $V_{\varphi}({\bm{s}})=\operatorname*{\mathbb{E}}_{{\bm{a}}\sim\pi(\cdot|{\bm{s}})}[Q_{\phi}({\bm{s}},{\bm{a}})]$ only conditioned on states is proposed to mitigate the out-of-distribution issue. It approaches the action distribution expectation of Q-value and can be optimized by expectile regression:
$\mathcal{L}_{V}(\varphi)=|\tau-\mathds{1}_{\{Q_{\phi}({\bm{s}}_{t}^{j},{\bm{a}}_{t}^{j})-V_{\varphi}({\bm{s}}_{t}^{j})<0\}}|(Q_{\phi}({\bm{s}}_{t}^{j},{\bm{a}}_{t}^{j})-V_{\varphi}({\bm{s}}_{t}^{j}))^{2},$ (6)
where $\tau\in(0,1)$ is a constant. Note that the action ${\bm{a}}$ will be sampled from the dataset according to the loss. Hence, we use rollouts in (5) to estimate returns by TD($\lambda$) (see Appendix.B). The objective can replace the TD-target that was originally used to optimize the Q-value. However, only updating the critic depending on model rollout is not enough, while the real state-action pairs from the offline dataset can supervise the critic to additionally learn in-distribution information. So the critic distribution is optimized on both model data and offline data:
$\mathcal{L}_{Q}^{1}(\phi)=-\frac{1}{HT}\sum_{t=1}^{T}\sum_{h=1}^{H}\log{\text{Pr}_{Q_{\phi}}(R_{t}^{h}|{\bm{s}}_{t}^{h})},\;\;\;\mathcal{L}_{Q}^{2}(\phi)=-\frac{1}{T}\sum_{t=1}^{T}\log{\text{Pr}_{Q_{\phi}}({\bm{r}}_{t}+\gamma V_{\varphi}({\bm{s}}_{t+1})|{\bm{s}}_{t})}$ (7)
where ${\bm{r}}_{t}$ is the reward from the offline dataset. Finally, the critic loss is the combination of the above.
$\mathcal{L}_{Q}(\phi)=\mathcal{L}_{Q}^{1}(\phi)+\mathcal{L}_{Q}^{2}(\phi)$ (8)
4.2 Model-based Policy Expansion for World Models Fine-tuning
After offline pretrain, we fine-tune the world model with online exploration to enhance performance and improve its generalization to similar tasks. However, directly using online algorithms suffers from non-recoverable performance drops potentially caused by the distribution shift and the change of the dynamic model. On the other hand, standard offline RL algorithms typically exhibit worse online fine-tuning performance due to their conservative design[20, 17]. Therefore, we try to bridge offline-online RL by fully utilizing offline pretraining policies[32].
Suppose that $\pi_{\beta}$ is the offline policy. Instead of directly fine-tuning it in the online stage, we optimize a new policy $\pi_{\psi}$ by gradually learning from prior policy $\pi_{\beta}$ and online exploring simultaneously. Concretely, given a current state ${\bm{s}}$, an action set is composed of actions generated by two policies: $\mathcal{A}=\{{\bm{a}}_{\beta}\sim\pi_{\beta}({\bm{s}}),{\bm{a}}_{\psi}\sim\pi_{\psi}({\bm{s}})\}.$ The actions are selected probabilistically based on their potential utilization value:
$P(\omega)[k]=\frac{\exp(Q_{\phi}({\bm{s}},{\bm{a}}_{k})/\alpha)}{\exp(Q_{\phi}({\bm{s}},{\bm{a}}_{\beta})/\alpha)+\exp(Q_{\phi}({\bm{s}},{\bm{a}}_{\psi})/\alpha)},\;\forall k\in\{\beta,\psi\},$ (9)
where $\alpha$ is temperature. According to this probability function, the two policies can be trained in a joint adaptive manner that has the advantage of adaptiveness in the sense of allowing generalizing of new abilities. In this way, the offline replay-buffer can be directly transferred to the online fine-tuning stage simplifying the complexity of balanced sampling[26].
In the model-based setting, the dynamic model also needs to be fine-tuned since the model primarily generates the policy update data. The world model benefits from its robust design in capturing temporal dependencies and modeling latent dynamics, which contributes to its stability during fine-tuning. However, the dynamic model might be influenced by the incorrect estimates learned from conservative offline training, making online fine-tuning more challenging. Thus, we develop an additional dynamic model for fine-tuning, expanding the policy to adapt to new tasks.
4.3 Reachability Estimation Function as Safety Guarantee
To ensure safety throughout the entire training process, prior works consider adopting the relaxation of the Lagrangian multiplier method on (4). The optimization problem is transformed into an unconstrained safe model-based problem with the Lagrangian multiplier $\lambda_{i}$:
$\max_{\pi_{\psi}}\min_{\lambda_{i}\geq 0}J^{\mathcal{R}}(\pi_{\psi})-\lambda_{i}(J^{\mathcal{C}}_{i}(\pi_{\psi})-d^{i}),\;\forall i\in\{1,\dots,C\}.$ (10)
However, the soft constraints will still lead to a certain chance of violation. It becomes more severe in the offline setting because of the difficulty in estimating the cost values. Hence, we replace soft constraints with hard constraints by the feasibility guidance approach. We first divide our optimization problem into two parts: the feasible part and the infeasible part.
For the feasible part, we want the agent to maintain safety within the feasible set and maximize the reward return. For the infeasible part, we want the policy to avoid the violations as little as possible and return to the feasible set. As the safe offline RL problem has another objective to optimize: $D(\pi_{\psi}||\pi_{b})$ ($\pi_{b}$ is the behavior policy), we can reformulate our offline safe model-based problem as follows:
[…]
where $V_{\varphi}^{r}({\bm{s}})=\operatorname*{\mathbb{E}}_{{\bm{a}}\sim\pi_{\psi}}[\sum_{t=0}^{\infty}\gamma^{t}r({\bm{s}}_{t},{\bm{a}}_{t})]$ and $V_{\varphi}^{c}({\bm{s}})=\operatorname*{\mathbb{E}}_{{\bm{a}}\sim\pi_{\psi}}[\sum_{t=0}^{\infty}\gamma^{t}c({\bm{s}}_{t},{\bm{a}}_{t})]$ represent the expected cumulative rewards and costs, $\mathcal{S}_{f}$ is the feasible set containing all feasible states. Note that the policy constraint is a soft constraint so the feasible part becomes a problem incorporating coupled hard and soft constraints. Thus, we use the reachability estimation function $u^{\pi}({\bm{s}})=\operatorname*{\mathbb{E}}_{\tau\sim\pi}[\max_{{\bm{s}}_{t}\in\tau}\mathds{1}_{({\bm{s}}_{t}|{\bm{s}}_{0},\pi)\in\mathcal{S}_{f}}]$ to handle the mixed constraints while prioritizing the hard constraints from RESPO[51]. It also assists the agent in reentering the feasible region even when it is in the infeasible region.
Therefore, we couple the out-of-distribution actions avoidance into the problem by using advantage functions. Also, we take advantage of the Augmented Lagrangian with a proximal relaxation method that is quite effective in safe model-based RL algorithms. The final problem is as follows:
$\max_{\pi_{\psi}}\;\operatorname*{\mathbb{E}}_{{\bm{s}}}\Big{[}(A^{r}({\bm{s}},{\bm{a}})-\Phi(V_{\varphi}^{c}({\bm{s}}),\lambda_{p}^{k},\mu^{k}))\cdot u^{\pi}_{\psi}-A^{c}({\bm{s}},{\bm{a}})\cdot(1-u^{\pi}_{\psi})\Big{]},\;\;\text{s.t.}\;\;D(\pi_{\psi}||\pi_{b})\leq\epsilon.$ (13)
where $A^{r}({\bm{s}},{\bm{a}})=Q_{\phi}^{r}({\bm{s}},{\bm{a}})-V_{\varphi}^{r}({\bm{s}})$ and $A^{c}({\bm{s}},{\bm{a}})=Q_{\phi}^{c}({\bm{s}},{\bm{a}})-V_{\varphi}^{c}({\bm{s}})$. We can finally obtain a closed-form solution of $\pi_{\psi}$, which we can use to extract the optimal policy. (For more details see Appendix.A.1)
$\pi^{*}({\bm{a}}|{\bm{s}})=\frac{1}{Z}w\cdot\pi_{b}({\bm{a}}|{\bm{s}}),$ (14)
$\mathcal{L}_{\pi}(\psi)=-\operatorname*{\mathbb{E}}_{\tau\sim\mathcal{D}}\Big{[}\sum_{t=1}^{T}(w\cdot\log\pi_{\psi}({\bm{a}}_{t}|{\bm{s}}_{t})+\eta H[\pi_{\psi}({\bm{a}}_{t}|{\bm{s}}_{t})])-\Phi(V_{\varphi}^{c}({\bm{s}}),\lambda_{p}^{k},\mu^{k})\cdot u^{\pi}_{\psi}\Big{]},$ (15)
$\text{where}\;\;\;w=u^{\pi}_{\psi}\cdot\exp(\beta_{1}A^{r}({\bm{s}},{\bm{a}}))+(1-u^{\pi}_{\psi})\cdot\exp(-\beta_{2}A^{c}({\bm{s}},{\bm{a}})).$ (16)
In the online fine-tuning stage, we apply policy expansion as follows: use $\pi_{b}$ as the frozen offline policy trained by (15) and learn a new policy $\pi_{\psi}$ without policy behavior constraint in (13). Thus, the online policy loss can change the advantage weight as:
$w=u^{\pi}_{\psi}\beta_{1}A^{r}({\bm{s}},{\bm{a}})-(1-u^{\pi}_{\psi})\beta_{2}A^{c}({\bm{s}},{\bm{a}}).$ (17)
For cost critic $Q^{c}(s,a)$ and value $V^{c}(s)$ training, we use the same way as Section 4.1. The whole algorithm training outline is shown in Appendix.B.
Figure 2: Offline experimental results. Comparing FOSP to baselines across five image-based safety tasks. The results for all three algorithms are obtained after training for 1 million steps. Reward: averaged episode reward return. Cost: averaged episode cost return. Cost Regret: averaged cost value throughout the training phase.
5 Experimental Results
In this section, we evaluate our method on different agents and tasks from the Safety-Gymnasium in simulation. We deploy it in motion planning control tasks on a real Franka robot as well. We aim to address the following issues: (1) How does FOSP perform in offline and online fine-tuning? (2) How does each component in FOSP contribute to its overall functionality? (3) Does FOSP successfully handle unseen safety regions in the real world?
5.1 Experimental Setup
Simulation Tasks
We consider five tasks on Safety-Gymnasium benchmark[11] environments. The agents need to navigate to predetermined goals without collision with hazards and vases. We measure the performance with three metrics: average undiscounted episode reward return (Reward), average undiscounted episode cost return (Cost) and average cost value throughout the training phase (Cost Regret). To evaluate the advantage of world models in handling high-dimensional features, we use 64$\times$64 pixels RGB image as inputs obtained from the first-person perspective of the agent, which is more representative of real-world environments. The offline datasets are sampled by three different behaviors: unsafe policy, safe policy and random policy to evaluate the performance of FOSP on diverse behavior samples. They are mixed in a 1:1:1 ratio, with each part containing 200 trajectories. For more information see Appendix.C.
Baselines
Because our work is the first approach in the safe model-based offline-to-online area to the best of our knowledge, there are not many baselines for us to compare. Prior works in offline safe RL use the model-free method which typically performs poorly in vision-only tasks. Therefore, we compare FOSP with the following two model-based baselines: SafeDreamer[8], an online powerful algorithm in safe model-based reinforcement learning without online planning and DreamerV3[5], a strong baseline in visual tasks but overlooks constraint violations.
Figure 3: Ablation studies. We evaluate ablations in SafetyPointGoal2 with means of three seeds. The vertical line divides the offline and online phases. (1M steps for offline and 0.75M steps for online)
Real Robot
The real robot environment includes a 7-DOF Franka Emika Panda robot as the controlled agent and a static third-person Intel RealSense camera to obtain images as inputs. Also, the agent has additional inputs: the robot end-effector pose information. We design the task SafeReach that controls the robot to reach a desired goal while avoiding collisions with predefined obstacles in its field of view. The rewards and costs are sparse and detected manually determining whether the robot reaches the goal or violates the constraints. The dataset consists of demonstration data, safety violation data, and failure data. Furthermore, we design some unseen safe regions and different targets that are not contained in the offline dataset. These unseen tasks are related to some undefined obstacles that need to be identified through observation which can not be solved by motion planning methods. The task further illustrates the practicality of our approach.
5.2 Main Results
Offline Pretrain
We show our offline training result in Figure 2. FOSP outperformed the similar safe algorithm SafeDreamer, excelling in decreasing costs and better balancing task performance with constraint avoidance. Furthermore, It has comparable or slightly lower performance than DreamerV3 while achieving nearly zero violations. Note that all methods struggle to get higher rewards in SafetyPointPush because the line-of-sight obstructions necessitate online exploration.
Online Fine-tuning
The results of online fine-tuning are shown on Table 1. It proves that our algorithm not only performs well offline but can also be fine-tuned to achieve better performance. Meanwhile, FOSP further reduces the costs during the online fine-tuning. We also compare our method with the OSRP version of SafeDreamer which can only be trained online. We notice that the SafeDreamer training exclusively online outperforms the offline-to-online fine-tuning approach. This may be because direct fine-tuning is affected by distribution shifts, making it difficult to improve performance in limited steps. So FOSP is necessary for us to bridge the safety of offline and online while maintaining its performance. (For more detail see Appendix.F)
Ablation Studies
In this section, we compare the effects of different components of the model on its performance in Figure 3. Specifically, we provide the following ablations: (i) learning value function without in-sample optimization (Eq.6) in Section 4.1; (ii) directly fine-tuning the model without model-based policy expansion (Eq.9) in Section 4.2; (iii) update safe actor without reachability estimation function in Section 4.3; (iv) design the penalty without the Augmented Lagrangian and use the feasible value function[52]. The results show that each component of our method individually contributes to ensuring offline training performance, further improvement during online fine-tuning, and maintaining near-zero cost constraint violations.
Table 1: Online fine-tuning results. Reward return and cost return of methods in offline-to-online fine-tuning (1M steps for offline and 0.5M steps for online fine-tuning). And results of online algorithms only trained online (1.5M steps). Means of three seeds.
Offline-to-Online Online-only
DreamerV3 SafeDreamer FOSP(Ours) SafeDreamer(+planning)
Tasks Reward Cost Reward Cost Reward Cost Reward Cost
PointGoal1 17.4$\rightarrow$25.8 82.3$\rightarrow$85.1 12.7$\rightarrow$19.2 1.5$\rightarrow$0.02 18.7$\rightarrow$21.5 1.4$\rightarrow$0.2 20.1 0.6
PointButton1 15.1$\rightarrow$20.8 167.7$\rightarrow$157.9 8.9$\rightarrow$10.7 10.2$\rightarrow$4.5 14.7$\rightarrow$18.1 9.5$\rightarrow$2.1 12.4 5.5
PointPush1 2.3$\rightarrow$14.6 30.9$\rightarrow$26.3 1.3$\rightarrow$10.3 17.8$\rightarrow$0.12 4.0$\rightarrow$13.2 18.1$\rightarrow$0.1 8.1 0.7
PointGoal2 9.5$\rightarrow$18.9 367.1$\rightarrow$290.2 3.6$\rightarrow$12.7 8.6$\rightarrow$0.3 8.1$\rightarrow$13.5 7.5$\rightarrow$0.23 13.0 1.7
CarGoal2 16.7$\rightarrow$24.1 231.0$\rightarrow$287.9 6.9$\rightarrow$9.8 0.7$\rightarrow$0.09 10.1$\rightarrow$14.5 1.6$\rightarrow$0.07 11.2 0.34
Average 20.8 171.1 12.5 1.01 16.6 0.54 12.9 1.77
5.3 Unseen Safe Tasks
In the real-world environment, robots will incur many safety constraints while performing various tasks. Our key challenge is how to safely train the robots to adapt to unseen constraints by safe learning-based methods. Following this insight, we introduce obstacles and targets of different shapes and sizes in Reach tasks, where the robot has only encountered a limited number of these during offline training. Our goal is to make the robot use a few trials to learn planning strategies in environments with new obstacles as safely as possible. Thus, we design three real-world transfer tasks (see details in Appendix.C) and show their results in Table 2.
Table 2: Real-world unseen tasks. We record the success rate (SR, %) and the constraint violation rate (CV, %) over 20 tests in three tasks while it will be labeled as a violation if it collides with an obstacle. The robot has fine-tuned 40 steps. See Appendix.C for more details.
Before FT During FT After FT
Tasks SR CV SR CV SR CV
1 40 50 45 40 50 30
2 30 50 35 45 35 40
3 20 60 40 40 50 30
In brief, FOSP can successfully finish these tasks safely, demonstrating its strong offline learning and adjustment capabilities. In contrast, only learning from the offline dataset makes it difficult to generalize to new scenarios. Its strong adaptability in fine-tuning showcases the robustness and great potential of FOSP.
6 Conclusion
We introduce FOSP, a novel method that integrates model-based reinforcement learning with offline-to-online fine-tuning to enhance safety in robotic tasks. FOSP addresses the challenges of balancing performance and safety during both offline and online phases using the reachability estimation function to unify different constraints and offline safe policies to guide online exploration. Our experimental results demonstrate that FOSP achieves robust performance in various safe planning tasks, especially those with high-dimensional vision inputs. Furthermore, we evaluate FOSP on unseen safety-critical tasks in dynamic real-world environments highlighting its practical value in safe few-shot fine-tuning. FOSP is the first approach to extend MBRL into safe offline-to-online training, showing promising results in the Safety-Gymnasium benchmark. Overall, FOSP shows a significant step forward in ensuring safe and efficient decision-making in embodied intelligence.
Appendix A Theoretical Interpretations
A.1 Derivation of the Final Optimization Problem
This section analyzes the derivation from Eq.(21) to Eq.(13) by using the Augmented Lagrangian and advantage function. First, we introduce Proposition 1.
Proposition 1.
The optimization objective of Eq.(11) is the necessary condition of $\max_{\pi}\operatorname*{\mathbb{E}}_{{\bm{s}}}[A^{r}({\bm{s}},{\bm{a}})\cdot\mathds{1}_{s\in\mathcal{S}_{f}}]$.
Proof. We start with the advantage function estimation in [53] and [23]. Suppose that $d_{\pi}({\bm{s}})=\sum_{t=0}^{\infty}\gamma^{t}p({\bm{s}}_{t}={\bm{s}}|\pi)$ is discounted state distribution with $\pi$, and $p({\bm{s}}_{t}={\bm{s}}|\pi)$ is the probability of the state ${\bm{s}}$ guided by $\pi$ for $t$ steps. The discounted sum of advantage expectation under $d_{\pi}({\bm{s}})$ can be represented as follows:
$\operatorname*{\mathbb{E}}_{{\bm{s}}_{1:\infty}\sim d_{\pi}({\bm{s}})}\left[\sum_{t=0}^{\infty}\gamma^{t}A_{\mu}({\bm{s}}_{t},{\bm{a}}_{t})\right]$
$=\operatorname*{\mathbb{E}}_{{\bm{s}}_{1:\infty}\sim d_{\pi}({\bm{s}})}\left[\sum_{t=0}^{\infty}\gamma^{t}\left(r({\bm{s}}_{t},{\bm{a}}_{t})+\gamma V_{\mu}({\bm{s}}_{t+1})-V_{\mu}({\bm{s}}_{t})\right)\right]$
$=\operatorname*{\mathbb{E}}_{{\bm{s}}_{1:\infty}\sim d_{\pi}({\bm{s}})}\left[-V_{\mu}({\bm{s}}_{0})+\sum_{t=0}^{\infty}\gamma^{t}r({\bm{s}}_{t},{\bm{a}}_{t})\right]$
$=-\operatorname*{\mathbb{E}}_{{\bm{s}}_{0}\sim d_{\pi}({\bm{s}}_{0})}\left[V_{\mu}({\bm{s}}_{0})\right]+\operatorname*{\mathbb{E}}_{{\bm{s}}_{1:\infty}\sim d_{\pi}({\bm{s}})}\left[\sum_{t=0}^{\infty}\gamma^{t}r({\bm{s}}_{t},{\bm{a}}_{t})\right]$
$=-J(\mu)+J(\pi),$
where $\mu$ is the behavior policy used to sample data. So we get the objective representation $J(\pi)=\operatorname*{\mathbb{E}}_{{\bm{s}}\sim d_{\pi}({\bm{s}})}\Big{[}\operatorname*{\mathbb{E}}_{{\bm{a}}\sim\pi({\bm{a}}|{\bm{s}})}[A_{\mu}({\bm{s}}_{t},{\bm{a}}_{t})]\Big{]}+J(\mu)$. Then, we have the following derivation:
$\max_{\pi}\operatorname*{\mathbb{E}}_{{\bm{s}}\sim d_{\pi}({\bm{s}})}[A^{r}_{\pi}({\bm{s}},{\bm{a}})]$ $\Rightarrow\max_{\pi}\operatorname*{\mathbb{E}}_{{\bm{s}}\sim d_{\pi}({\bm{s}})}\Big{[}\operatorname*{\mathbb{E}}_{{\bm{a}}\sim\pi({\bm{a}}|{\bm{s}})}[A_{\mu}^{r}({\bm{s}}_{t},{\bm{a}}_{t})\Big{]}$
$=\max_{\pi}\operatorname*{\mathbb{E}}_{{\bm{s}}\sim d_{\pi}({\bm{s}})}\Big{[}\operatorname*{\mathbb{E}}_{{\bm{a}}\sim\pi({\bm{a}}|{\bm{s}})}[A_{\mu}^{r}({\bm{s}}_{t},{\bm{a}}_{t})\Big{]}+J(\mu)$
$=\max_{\pi}\;J(\pi)$
$=\max_{\pi}\;\operatorname*{\mathbb{E}}[V_{\pi}^{r}({\bm{s}})].$
The last equation holds by the definition of $V_{\pi}^{r}({\bm{s}})$. $\square$
Due to the symmetry of $\max_{\pi}\operatorname*{\mathbb{E}}[V_{\pi}^{r}({\bm{s}})]$ and $\max_{\pi}\operatorname*{\mathbb{E}}[V_{\pi}^{c}({\bm{s}})]$, the same proof also applies to $\max_{\pi}\operatorname*{\mathbb{E}}[V_{\pi}^{c}({\bm{s}})]$. So we have proposition 2.
Proposition 2.
The optimization objective of Eq.(12) is the necessary condition of $\max_{\pi}\operatorname*{\mathbb{E}}_{{\bm{s}}}[-A^{c}({\bm{s}},{\bm{a}})\cdot\mathds{1}_{s\notin\mathcal{S}_{f}}]$.
The problem converts to:
[…]
However, it is hard for us to get a close form of $\pi_{\psi}$ in this problem. Therefore, we will first disregard the constraints of the behavior policy and address the hard safety constraints. The Augmented Lagrangian is used to deal with the feasible part. Thus, we use the following relaxation:
$\max_{\pi_{\psi}}\min_{\lambda\geq 0}\left[A^{r}({\bm{s}},{\bm{a}})\mathds{1}_{s\in\mathcal{S}_{f}}-\lambda V_{\varphi}^{c}({\bm{s}})+\frac{1}{\mu_{k}}(\lambda-\lambda_{k})^{2}\right],$ (20)
where $\mu_{k}$ is a non-decreasing penalty term corresponding to gradient step $k$. By smoothing the left-hand side term in Eq.(18), the last term ensures that $\lambda$ stays near its previous estimate. Differentiating Eq.(18) with respect to $\lambda_{p}^{k}$ leads to the following update rule for the Lagrange multiplier:
$\lambda_{p}^{k+1}=\begin{cases}\lambda_{p}^{k}+\mu^{k}V_{\varphi}^{c}&\text{ if }\lambda_{p}^{k}+\mu^{k}V_{\varphi}^{c}\geq 0,\\ 0&\text{ otherwise }.\end{cases}$
The objective becomes the following form:
$J(\pi,\lambda_{k},\mu_{k})=A^{r}({\bm{s}},{\bm{a}})\mathds{1}_{s\in\mathcal{S}_{f}}-\Psi\left(V_{\varphi}^{c},\lambda_{p}^{k},\mu^{k}\right),$
$\Psi\left(V_{\varphi}^{c},\lambda_{p}^{k},\mu^{k}\right)=\begin{cases}\lambda_{p}^{k}V_{\varphi}^{c}+\frac{\mu^{k}}{2}(V_{\varphi}^{c})^{2}&\text{ if }\lambda_{p}^{k}+\mu^{k}V_{\varphi}^{c}\geq 0,\\ -\frac{\left(\lambda_{p}^{k}\right)^{2}}{2\mu^{k}}&\text{ otherwise }.\end{cases}$
Then, we utilize the REF function to connect the two parts of the problem. The REF function is defined as follows:
Definition 1.
The reachability estimation function (REF) $u:\mathcal{S}\rightarrow[0,1]$ is:
$u^{\pi}({\bm{s}}):=\operatorname*{\mathbb{E}}_{\tau\sim\pi}[\max_{{\bm{s}}_{t}\in\tau}\mathds{1}_{({\bm{s}}_{t}|{\bm{s}}_{0},\pi)\in\mathcal{S}_{f}}]$
And the problem becomes:
$\max_{\pi_{\psi}}\;\operatorname*{\mathbb{E}}_{{\bm{s}}}\Big{[}V_{\varphi}^{r}({\bm{s}})\cdot u^{\pi}_{\psi}-V_{\varphi}^{c}({\bm{s}})\cdot(1-u^{\pi}_{\psi})\Big{]},\;\;\text{s.t.}\;\;V_{\varphi}^{c}({\bm{s}})=0,\;\;D(\pi_{\psi}||\pi_{b})\leq\epsilon.$ (21)
We introduce a proposition from [10] to solve the soft constraints from offline training to provide a solution of Eq.(13).
Proposition 3.
Suppose that the optimizing problem has the following form:
$\max_{\pi}\operatorname*{\mathbb{E}}_{{\bm{a}}\sim\pi}[A(s,a)]\;\;\text{s.t.}\;D(\pi_{\psi}||\pi_{b})\leq\epsilon.$
The optimal solution will satisfy:
$\pi^{*}({\bm{a}}|{\bm{s}})\;\propto\;\exp(\alpha A(s,a))\pi_{b}({\bm{a}}|{\bm{s}})$
Proof. The Lagrange function can be formulated as follows:
$L(\pi,\lambda,\mu)=\operatorname*{\mathbb{E}}_{{\bm{a}}\sim\pi}[A({\bm{s}},{\bm{a}})]-\lambda(D(\pi_{\psi}||\pi_{b})-\epsilon),$
Then we take the partial derivative with respect to $\pi$ and set it to 0:
$\frac{\partial L}{\partial\pi}=A(s,a)-\lambda\log\pi_{b}({\bm{a}}|{\bm{s}})+\lambda\log\pi({\bm{a}}|{\bm{s}})+\lambda=0$
So we have:
$\pi^{*}({\bm{a}}|{\bm{s}})=\frac{1}{Z}\exp(\alpha A(s,a))\pi_{b}({\bm{a}}|{\bm{s}}),$
where $Z$ is a normalizing constant. $\square$
Therefore, we have a solution to our optimization objective (13). The solved $\pi_{\psi}$ can be explicitly expressed by Eq.(15).
Appendix B Implementation Details
The TD($\lambda$) return can be estimated as follows:
$R_{t}^{H}=V_{\varphi}({\bm{s}}_{t}^{H}),\;R_{t}^{h}={\bm{r}}_{t}^{h}+\gamma((1-\lambda)V_{\varphi}({\bm{s}}_{t}^{h})+\lambda R_{t}^{h+1})$ (22)
where $H$ is the horizon of the model and $h\leq H$.
To implement the reachability estimation function $u_{\xi}$, we use the theorem introduced by [51] (Theorem 1). The REF function can be trained as follows:
$u({\bm{s}}_{t})=\max\{\mathds{1}_{s\in\mathcal{S}_{f}},\gamma_{u}u({\bm{s}}_{t+1})\},$ (23)
where $\gamma_{u}$ is a discount parameter $10<<\gamma_{u}<$ 1 to ensure convergence of $u({\bm{s}})$. Meanwhile, we ensure that its learning rate is greater than that of the critic and policy, but less than that of the Lagrangian multiplier updates.
Algorithm 1 FOSP: Fine-tuning Offline Safe Policy through World Models
0:  Offline dataset $\mathcal{D}$, policy ${\pi}_{\psi}$, critics $Q_{\phi_{1}}$ and cost critics $Q^{c}_{\phi_{2}}$, value network $V_{\varphi_{1}},V^{c}_{\varphi_{2}}$, reachability function network $u_{\xi}$, world model $p_{\theta}$, policy rollout length $H$, number of offline training steps $N_{\text{offline}}$, number of online fine-tuning steps $N_{\text{online}}$.
1:  Initialize neural network parameters $\theta,\psi,\phi_{i},\varphi_{i},\xi$.$(i=1,2)$
2:  for $i=1,2,3,\cdots,N_{\text{offline}}$ do
3:     Sample a batch of trajectories $B\sim\mathcal{D}$.
4:     Compute model state ${\bm{s}}_{t}\sim p_{\theta}({\bm{s}}_{t}|{\bm{s}}_{t-1})$
5:     Update $\theta$ using Eq. (D).
6:     Generate $H$-step latent policy rollouts using world model $p_{\theta}$.
7:     Compute target estimation $R_{t}$ by TD($\lambda$).
8:     Update $V_{\varphi_{1}},V_{\varphi_{2}}$ using Eq. (6).
9:     Update $Q_{\phi_{1}},Q_{\phi_{2}}$ using Eq. (8).
10:     Update ${\pi}_{\psi}$ using Eq. (15).
11:     Update $u_{\xi}$ using Eq. (23).
12:  Initialize a new policy $\pi_{\psi^{\prime}}$ and freeze offline policy $\pi_{\psi}$ , transfer critics and values.
13:  for $i=1,2,3,\cdots,N_{\text{online}}$ do
14:     Switch a policy $\pi$ according to Eq.(9) and rollout the policy in the environment for an episode to collect a new trajectory $\tau$.
15:     $\mathcal{D}=\mathcal{D}\cup\tau$.
16:     for each training step do
17:        Sample a batch of trajectories from $\mathcal{D}$.
18:        Update $\theta$ using Eq. (D).
19:        Generate $H$-step latent policy rollouts using world model $p_{\theta}$.
20:        Compute target estimation $R_{t}$ by TD($\lambda$).
21:        Update $V_{\varphi_{1}},V_{\varphi_{2}}$ using Eq. (6).
22:        Update $Q_{\phi_{1}},Q_{\phi_{2}}$ using Eq. (8).
23:        Update ${\pi}_{\psi^{\prime}}$ using Eq. (17).
24:        Update $u_{\xi}$ using Eq. (23).
Appendix C Experimental Details
C.1 Real World Experiments
The experiments verify the effectiveness of the algorithm by controlling a robotic arm to avoid obstacles and complete a reach task to the target objection, called SafeReach. As shown in Figure 4(c), the main hardware setup consists of three parts: a Franka Panda robotic arm, an experimental platform for placing obstacles and target objections, and an Intel RealSense D435 camera for perception. On the experimental platform, we used white lines to designate a 25cm×$\times$25cm operational area as the task space for the robotic arm.
SafeReach
For the task of reaching, the robot needs to consider the image information acquired by the camera along with the current joint posture information as planning conditions to determine the movement decision at the next step. As shown in Figure 4, the robotic arm needs to navigate within the designated movement space, bypassing the obstacles which are represented by the green rectangular blocks, to touch the predefined target which is represented by the red square block.
Figure 4: Simulation and real-world environment. Different tasks and agents in the simulation and real-world. (a) Simulation tasks: we consider four different tasks in the simulation: Push1, Goal1, Button1, Goal2. (from upper-left to lower-right) The preceding words indicate the task type and the following number represents the difficulty level. (b) Agent type: Car (upper) and Point (lower). (c) Real-robot setup: we use raw images as inputs and enable the robotic arm to complete obstacle avoidance tasks safely.
Figure 5: Unseen Tasks. We choose three unseen tasks to show the generalization of our algorithm with different obstacles and goals.
Dataset Collection
To train the model for the SafeReach task, 200 motion trajectories were collected as a dataset by teleoperating the robotic arm using a 3dconnexion SpaceMouce. The dataset consists of the following: 120 trajectories where the robotic arm did not touch any obstacles and successfully reached the target object; 30 trajectories where it did not touch any obstacles but also did not successfully reach the target; 40 trajectories where it touched obstacles but successfully reached the target; 10 trajectories where it touched obstacles and did not successfully reach the target. Each trajectory contains 80 sets of data points at different time steps. Each data point includes: state: a 64$\times$64 RGB image captured by the camera, a 6-dimensional pose list of the robotic arm; action: a 6-dimensional list recording the changes in the robotic arm’s pose; cost: 1-dimensional cost index indicating whether the robotic arm touched an obstacle; reward 1-dimensional reward index indicating whether the object reached the target object. The labeling methods for the cost and reward index are shown in Figure 6. The cost is set to 1 when the robotic arm collides with an obstacle and it is set to 0 otherwise. The reward is set to 1 when the object touches the target and it is set to 0 otherwise. We defined 3 placement points for target objects and 6 placement points for obstacles in the scenario. By restricting the placement positions of the target objects and obstacles, we ensure that the model can converge more efficiently on a limited dataset.
Figure 6: Dataset collection. Each row in the figure represents five frames of the robotic arm’s motion captured from the same single trajectory. From the moment the robotic arm starts touching an obstacle to the moment the obstacle’s posture stops changing, the cost label in the dataset is marked as 1, while in other cases it is marked as 0, as shown in the red-framed part in the image. When the robotic arm touches the target object, the reward label in the dataset is marked as 1, as shown in the yellow-framed part in the image.
Transfer Tasks
To validate the generalization of the algorithm, we design a series of transfer experiments. As illustrated in the left side of Figure 5, the original task involves 2 green rectangular obstacles and 1 red cylindrical target objection. We design three different transfer tasks, depicted on the right side of Figure 5. The tasks include altering the shape of the obstacles, changing the shape of the target object, and increasing the number of obstacles. These experiments aim to verify the algorithm’s generalization in avoiding obstacles of various shapes and quantities, as well as in reaching target objects of different shapes.
C.2 Simulation Environment
SafetyGoal
The agent’s objective is to reach a goal while avoiding obstacles in its environment. Each time the agent successfully reaches the designated goal, the environment randomly generates a new one. The agent earns rewards for approaching or reaching the goal but incurs penalties when it encounters obstacles. These obstacles include immovable hazards and movable vases. The agent’s observation state is represented by the images from its front and back. Two specific tasks are selected: the Point-Goal task and the Car-Goal task, which are performed by the Point and Car agents, respectively. The Point agent is a robot that operates on a 2D plane and is capable of rotating and moving both forward and backward. On the other hand, the Car agent, which is slightly more complex, features two independent parallel wheels and a freely rolling rear wheel.
SafetyButton
The agent’s goal is to navigate around both stationary and moving obstacles in the environment to press one of several target buttons. Similar to the goal task, the agent is rewarded for approaching or pressing the target button. The Button task introduces dynamic obstacles that move quickly along set paths, making it more challenging than the Goal task due to the costs incurred from collisions with these moving obstacles.
SafetyPush
The agent’s goal is to push an object to reach a desired goal while avoiding hazards and vases in the environment. It will get rewards when the object successfully pushes the yellow object to the goal. The agent does not fully control the object’s movement, making the task more difficult to handle. And there are also some line-of-sight obstruction problems in vision-only situations.
C.3 Simulation Dataset
We collect simulation data from three policies: random policy, safe policy and unsafe policy. Each policy collected 200 trajectories for each task, which were then combined into the final dataset. This approach ensures that our dataset uniformly includes all types of data: low cost with low return, low cost with moderate return, high cost with low return, and high cost with high return. Thus, it can make a better trade-off in training different offline policies. The safe policy is trained by SafeDreamer[8] and the unsafe policy is trained by DreamerV3[5] for 1.5M steps on each task. It shows that unsafe policies tend to collect data with higher returns but significant cost uncertainty in these scenarios.
Figure 7: The dataset distribution on five simulation tasks. We plot the reward return and the cost return of every data trajectory.
Appendix D Baselines
Figure 8: Online experimental results. Comparing FOSP to baselines across five image-based safety tasks in online fine-tuning. The results for all three algorithms are obtained after fine-tuning for 750,000 steps. As we can see, FOSP can maintain safety and achieve better performance during fine-tuning.
DreamerV3
DreamerV3[5] is a model-based reinforcement learning method that outperforms specialized methods across over 150 diverse tasks. It has a stable performance without adjusting hyperparameters in exploring farsighted strategies from pixels and sparse rewards in an open world. However, it overlooks the safety considerations in the environment, which brings high costs in safety-critical tasks.
It uses RSSM as the world model, and the loss function is similar to Eq.(D) but lacks the cost head.
$\mathcal{L}^{\text{model}}(\theta)=$ $\operatorname*{\mathbb{E}}_{\tau\sim\mathcal{D}}\Big{[}\sum_{t=1}^{T}-\ln p_{\theta}({\bm{x}}_{t}\mid{\bm{s}}_{t})-\ln p_{\theta}({\bm{r}}_{t}\mid{\bm{s}}_{t})+\mathbb{D}_{KL}[sg(q_{\theta}({\bm{z}}_{t}|{\bm{h}}_{t},{\bm{x}}_{t}))||p_{\theta}^{i_{t}}({\bm{z}}_{t}|{\bm{h}}_{t})]+$
$\beta\mathbb{D}_{KL}[q_{\theta}({\bm{z}}_{t}|{\bm{h}}_{t},{\bm{x}}_{t})||sg(p_{\theta}^{i_{t}}({\bm{z}}_{t}|{\bm{h}}_{t}))]\Big{]}.$
And a simple actor-critic framework is used in its decision-making part.
[…]
In the offline pretraining phase, we removed the online planning component and used it only for online fine-tuning. To achieve robust performance in the offline phase, we train a latent dynamics model ensemble and use the uncertainty estimation approach[14]:
$\hat{{\bm{r}}}_{\theta}({\bm{s}}_{t},{\bm{a}}_{t})={\bm{r}}_{\theta}({\bm{s}})-\alpha\cdot\text{std}(\{\log(p_{\theta}^{i}({\bm{z}}|{\bm{h}}))\}_{i=1}^{N}),$
where ${\bm{r}}_{\theta}({\bm{s}})$ is the reward predicted by the reward encoder.
SafeDreamer
SafeDreamer[8] incorporates Lagrangian-based methods into world model planning processes and achieves nearly zero-cost performance on various tasks. It performs well in high-dimensional vision-only input safety-critical tasks, surpassing the prior works[7, 40] and balancing performance and safety. The framework uses the same world models based on DreamerV3[5]. We use the BSRP-Lag version of the model, which utilizes the Lagrangian method in background safety-reward planning that avoids online planning. And this version of SafeDreamer also achieves the best performance of the three versions.
It adds a cost head into world models like our method and uses the same loss Eq.(D) in model training. It constructs a cost critic $V_{\phi}^{c}(R_{t}|{\bm{s}}_{t})$ like $V_{\phi}(R_{t}|{\bm{s}}_{t})$ and utilize Augmented Lagrangian method to update the actor:
$\mathcal{L}(\theta)=-\sum_{t=1}^{T}R^{\lambda}(s_{t})+\eta\mathrm{H}\left[\pi_{\theta}\left(a_{t}\mid s_{t}\right)\right]-\Psi\left(C^{\lambda}(s_{t}),\lambda_{p}^{k},\mu^{k}\right)$
$\Psi\left(C^{\lambda}(s_{t}),\lambda_{p}^{k},\mu^{k}\right),\lambda_{p}^{k+1}=\begin{cases}\lambda_{p}^{k}\Delta+\frac{\mu^{k}}{2}\Delta^{2},\lambda_{p}^{k}+\mu^{k}\Delta&\text{ if }\lambda_{p}^{k}+\mu^{k}\Delta\geq 0\\ -\frac{\left(\lambda_{p}^{k}\right)^{2}}{2\mu^{k}},0&\text{ otherwise }\end{cases}$
where $\delta=C^{\lambda}(s_{t})-b$.
Appendix E Hyperparameters
The experiments for FOSP were conducted in a Python 3.10 environment with JAX 0.4.26. Our setup included CUDA version 12.1, running on Ubuntu 20.04. The hardware used comprised of four GeForce RTX 4090 GPUs and an Intel(R) Xeon(R) Platinum 8358P CPU @ 2.60GHz CPU.
Table 3: Hyperparameters for FOSP
Module Name Symbol Value
World Model Number of latent $N_{l}$ 48
Classes per latent $C_{l}$ 48
Batch size $B$ 64
Batch length $T$ 16
Learning rate $l_{wm}$ $10^{-4}$
Coefficient of KL-divergence $\beta$ 0.1
Generation horizon $H$ 15
Augmented Lagrangian Penalty term $\nu$ $5^{-9}$
Initial Penalty multiplier $\mu^{0}$ $1^{-6}$
Initial Lagrangian multiplier $\lambda_{p}^{0}$ 0.01
Actor Critic Discount horizon $\gamma$ 0.997
Reward lambda $\lambda_{r}$ 0.95
Cost lambda $\lambda_{c}$ 0.95
Expectile $\tau$ $0.8$
AWR temperature $\beta_{1}$, $\beta_{2}$ 10
REF discount $\gamma_{u}$ 0.99
PEX temperature $\alpha$ 10
Actor entropy regularize $\eta$ $3\cdot 10^{-4}$
Learning rate $l_{ac}$ $3\cdot 10^{-5}$
General Number of MLP layers 5
Number of MLP layer units 512
Action repeat 4
Appendix F Additional Experimental Results
We present the training curves of FOSP during online fine-tuning in simulation experiments in Figure 8. The results show that with minimal fine-tuning, FOSP achieves better performance, outperforming SafeDreamer. Although it does not match the performance of DreamerV3, FOSP achieves nearly zero-cost of the agent throughout the entire process.
Appendix G Limitations and Future Work
Limitations
Despite achieving good results in many simulated environment experiments, the proposed framework still has some limitations. During the fine-tuning process with new trajectory data, the model struggles to update and converge efficiently with zero constraint violations. Also, we find that the model performs poorly in tasks with occluded vision. In experiments with real robots, our test results remain unstable and are sensitive to hyperparameter adjustments during training. This indicates the need for significant effort in tuning the model’s hyperparameters to achieve better performance. In transfer experiments, we find that when the positions of obstacles and target objects are random, the model has difficulty in converging with limited data despite fine-tuning. This is why we constrain the placement range of obstacles and target objects during our experiments, which also highlights the model’s high dependency on the dataset distribution.
Future Work
The possible future work will concentrate on three main fronts: improving safety considerations, assessing the algorithm across diverse tasks, and fortifying its robustness in real-world unseen tasks. To further decrease the violations, the safety component of our method can be improved during the offline-online stage. The SafeReach task is a preliminary experiment in the real world and it is possible to expand our method to general robot tasks such as grasping, pulling, and pushing. Moreover, the real-world task performance of our method can be further improved. To address the issue of inaccurate position determination from a single perspective and line of sight occlusion, the multi-view RL shows its efficiency in catching better features from images. What’s more, increasing the dataset capacity and uniforming the dataset distribution can probably enhance the robustness of the performance and expand it to more complex scenarios.